First of all would like to thank you for you help,I would like to know is this consider a complex equation or number and how would i solved it algebraically,
5-2/x-6=10-2x/x-6..
: First of all would like to thank you for you help,I would like to know is this consider a complex equation or number and how would i solved it algebraically,
5-2/x-6=10-2x/x-6..

Answer by ankor@dixie-net.com(4538)

is this considered a complex equation or number and how would i solved it algebraically?
:
It's not a complex equation, or a complex fraction
:
5-(2/(x-6)) = 10-(2x/(x-6))
:
The fractions have the same denominator, combine on the left, numbers on the right
(-2/(x-6)) + (2x/(x-6)) = 10 - 5
(2x-2)/(x-6) = 5
:
Multiply both sides by (x-6) to get rid of the denominator
2x - 2 = 5(x-6)
:
2x - 2 = 5x - 30
:
-2 + 30 = 5x - 2x
:
28 = 3x
x = 28/3; 91/3
:
:
Check solution using x = 9.33 on a calc
5-(2/((9.33-6))) = 10-2(9.33)/((9.33-6)))
5-(2/(3.33)) = 10-(18.66/(3.33))
5 - .6 = 10 - 5.6; confirms our solution